What is the awnser to this question?: l=12inches, w=4inches? Lenght and with. I have to find the perimeter and the area of the rectangle or square with the given demensions. Explain how u got it please.
Found 2 solutions by jim_thompson5910, nyc_function:
Answer by jim_thompson5910(35256)

Area:
A = LW

A = 12*4

A = 48

Area is 48 square inches.

-------------------------------------------------------
Perimeter

P = 2L + 2W

P = 2(12)+2(4)

P = 24+8

P = 32

Perimeter is 32 inches.


If you need more help, email me at jim_thompson5910@hotmail.com

Also, please consider visiting my website: http://www.freewebs.com/jimthompson5910/home.html and making a donation. Thank you

Jim

Answer by nyc_function(2741)

Use the following two formulas:
For area use: A = Length times width
For perimeter use: P = 2(length) + 2(width)

L = 12, W = 4

A = 12 times 4

A = 48in^2

The area is 48 squared inches.

==============================
P = 2(12) + 2(4)

P = 24 + 8

P = 32 inches

The perimeter is 32 inches.

The figure given is a rectangle not a square.